(y+1)^2=8(x-4)
How do I identify the lines of symmetry?

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

In the general equation (y-k)^2 = a(x-h), the line of symmetry is a horizontal line and it is the equation y = k

In this case, the value of k is k = -1, so the answer is y = -1 (this is the horizontal line at y = -1)

Answer by stanbon(75887)

(y+1)^2=8(x-4)
How do I identify the lines of symmetry?
-------
Because the equation is basically x = y^2 the graph
would be a parabola opening to the right from
vertex (4,-1)
------
The line of symmetry is a horizontal line passing
thru (4,-1):: y = -1 is the line of symmetry.
----------
Cheers,
Stan H.